OpenStudy (tamara7301):

s=2πrh solve for the variable r

OpenStudy (anonymous):

What does it mean to solve for a variable? It means to perform operations on the equation in order to isolate the variable asked for on the left side of the = sign and all else on the right side of the = sign 1) S=2πrh, for variable r Since all the terms are multiplied we divide both sides by those variables which will be moved to the other side in order to get r by itself diveid by 2πh S/2πh = r And that's all there's to it And to make it more elegant just flip the eqaution to r = S/2πh 2) V= πr^2h, for variable h Same procedure in this one Divide both sides byπr^2 V/πr^2 = h h = V/πr^2

OpenStudy (anonymous):

Hi, Um you are in history! you should be in math. But any way the answer is up top.